Question 2. Based on market research, a film production company in Ectenia obtains the following information about the demand and production costs of its new DVD: Demand: P = 1,000 − 10Q Total Revenue: TR = 1,000Q − 10Q2 Marginal Revenue: MR = 1,000 − 20Q Marginal Cost: MC = 100 + 10Q where Q indicates the number of copies sold and P is the price in Ectenian dollars. a. Find the price and quantity that maximize the company’s profit. (Monopoly profit) b. Find the price and quantity that would maximize social welfare. (Competitive profit


1
Expert's answer
2021-12-27T08:59:13-0500

a. The price and quantity that maximize the company’s profit are:

MR = MC,

1,000 − 20Q = 100 + 10Q,

30Q = 900,

Q = 30 units,

P = 1,000 - 10×30 = 700.

b. The price and quantity that would maximize social welfare are:

P = MC,

1,000 − 10Q = 100 + 10Q,

20Q = 900,

Q = 45 units,

P = 1,000 - 10×45 = 550.